# Semantics-aware Test-time Adaptation for 3D Human Pose Estimation

### v1

@@ 1 Introduction
This paper tackles the adaptation of 3D human pose estimation models to in-the-wild videos. A practical alternative to improve predictions is Test-Time Adaptation (TTA). TTA directly uses the (unlabeled) target sequences to fine-tune the pre-trained model during inference. Often, pre-trained models exhibit a misalignment between their projected keypoints and 2D image evidence. Based on this result, previous TTA methods (Guan et al., 2022; Nam et al., 2023) usually penalize the distance between the 2D-projections and 2D evidence provided by either ground truth or estimated 2D poses. However, such a scheme has two main challenges: 1) Depth ambiguity, which arises because numerous 3D pose solutions correspond to the same 2D pose. This ambiguity may improve 2D alignment but worsen 3D accuracy and plausibility. (2) Missing 2D keypoints, e.g., under occlusions and truncations. These keypoints have no guidance for improvement.
@@ 2.1 Test-time Adaptation in 3D Pose Estimation
However, ground truth 2D poses are not always available in many real-world scenarios. Similar to our method, CycleAdapt (Nam et al., 2023) operates on estimated 2D data and introduces a cyclic framework, where the motion denoising module smooths 3D pseudo labels to mitigate the impact of noisy 2D inputs. Nonetheless, the inherent 2D-to-3D ambiguity still significantly degrades model performance, and unguided predictions often occur when 2D inputs are unreliable. Our method addresses this by leveraging explicit semantic understanding, i.e., , actions, to guide the model in predicting specific motion patterns. This is the first approach to emphasize high-level semantic information to reduce the 2D-to-3D solution space in test-time 3D human pose estimation.
@@ 2.2 Human Motion Priors
Leveraging motion prior as supervision is commonly applied in video-based human pose estimation due to the lack of 3D annotations with RGB input. Trained on a large and diverse human motion database AMASS (Mahmood et al., 2019; Ionescu et al., 2013), the human prior provides a regularized motion space. Previous works mainly drag the predicted or generated motion to the regularized motion space to achieve the refinement (Kocabas et al., 2020; Shin et al., 2024; Rempe et al., 2021), infilling (Yuan et al., 2022; Zhang et al., 2021), denoising (Nam et al., 2023) goals. Specifically, VIBE (Kocabas et al., 2020) extended VAE (Pavlakos et al., 2019) and adversarial (Kanazawa et al., 2018) feasible prior from pose into motion space to improve temporal continuity of movement. GLAMR (Yuan et al., 2022) learned a VAE motion infiller to infill motion where the moving human is occluded. CycleAdapt utilizes the prior to denoise 3D meshes and improves human pose estimation during test-time with noisy 2D poses. However, those applications of human motion prior only optimize the predicted meshes of the segment to be valid motion without consideration of the motion semantics. Compared to other works, we are the first to split regularized motion space according to text labels extracted from the video segment to benefit motion prediction.
Figure 2: Overall framework: we perform model fine-tuning and 2D pose update in each epoch. Model Fine-tuning (Left)-The extracted video segment and its text label will be used to update $f^{\text{HPE}}$ with 2D projection, alignment, and smooth losses. 2D Pose Update (Right)- After model fine-tuning in each epoch, all images in the current test video are processed through $f^{\text{HPE}}$ to obtain predicted 2D and update the 2D keypoints using EMA. Subsequently, all video segments are checked to fill in any missing 2D keypoints by verifying conditions of motion-text alignment.
@@ 3.1 Problem Setup
3.2 Motion-language Models
A motion-language model integrates human motion and natural language modality. It utilizes pre-trained language models (Radford, 2018; Kenton & Toutanova, 2019), to incorporate rich semantic space for improved motion representation. This is achieved by training with motion-text data, which enables tasks like text-conditioned motion generation (Zhang et al., 2022a, 2023) and motion captioning (Guo et al., 2022; Jiang et al., 2023). Our focus is on leveraging motion-language model to guide motion semantics for unlabeled test videos in TTA. We choose MotionCLIP (Tevet et al., 2022) for its text-aligned motion representation, which is optimally suited to our task.
@@ 4.1 Semantics-aware Motion Prior
More details about the VLM prompts and label verification can be found in Supplementary Materials.
Video segment processing. After the text label for motion in each image is obtained, we aim to align the the projected motion and text features within the shared space of MotionCLIP. To process video segments for training, we divide the video with a stride of 1, and only keep the segments where all images share the same text label. This is to avoid misalignment caused by mixed motions as much as possible. To further improve alignment, we adopt weighted sampling and prioritize those segments whose motion predictions deviate from the true semantics. Specifically, the sampling weight is defined as 1 minus cosine similarity between predicted motion and its text label in MotionCLIP.
@@ 4.3 Overall Training
${\mathcal{L}}_{\textnormal{smooth}}=||\bm{\hat{\beta}}-\bm{\bar{\beta}}||_{1}+||\bm{\hat{\theta}}-\bm{\bar{\theta}}||_{1},~{}$ (9)
where the $\bm{\bar{\beta}}$ comes from the motion denoise module, and $\bm{\bar{\theta}}$ is the averaged shape estimation with a sliding window operation. Together with motion-text feature alignment loss, 2D projection loss, and loss weight $\lambda$, our training objective in each epoch can be summarized as:
@@ 5 Experiments
Table 2: Test-time adaptation comparison results on EgoBody (Zhang et al., 2022b) and 3DPW (Von Marcard et al., 2018) based on HMR2.0 network (Goel et al., 2023). Our method achieves better performance than CycleAdapt (Nam et al., 2023).
@@ 5.1 Setup
Datasets. We follow the adaptation tasks from previous work (Zhang et al., 2020; Nam et al., 2023), using Human3.6M (Ionescu et al., 2013) as the labeled training dataset and 3DPW (Von Marcard et al., 2018) and 3DHP (Mehta et al., 2017) as the unlabeled test datasets. Human3.6M is a widely used indoor dataset comprising 3.6 million images annotated with 2D and 3D labels. Despite its extensive pose diversity, the HMR network (Kanazawa et al., 2018) pre-trained on this dataset often struggle to generalize to in-the-wild settings due to its limited variation in appearance and environmental factors. The 3DPW test set comprises 37 videos and presents a significant challenge as it is an outdoor dataset featuring diverse scenarios, including subject occlusion and varied environments. The 3DHP test set includes 6 videos with both indoor and outdoor scenarios, and it presents an additional challenge by featuring some poses not seen in Human3.6M. Furthermore, we validate our method on a egocentric dataset EgoBody (Zhang et al., 2022b), which presents severe body truncation with missing 2D evidence.
@@ 5.2 Implementation Details
At the start of test-time adaptation for each test video, the model parameters are initialized with the pre-trained values, following (Nam et al., 2023). We employed the Adam optimizer (Kingma & Ba, 2014) with parameters set to $beta1$ = 0.5, $beta2$ = 0.9, and a learning rate of 5.0e-5. A cosine scheduler is used with a minimum learning rate of 1.0e-6. The input images are resized to 224$\times$224, and the frame number of each video segment is 60. We use a batch size of 4 and the total training epoch is 6. The hyperparameters are $\lambda_{1}=0.1$, $\lambda_{2}=0.2$, $\sigma=0.75$, $\alpha=0.9$. We use Openpose (Cao et al., 2017) to provide 2D poses. The comparison methods include state-of-the-art test-time adaptation methods: BOA (Guan et al., 2021), DynaBOA (Guan et al., 2022), DAPA (Weng et al., 2022) and CycleAdapt (Nam et al., 2023).
Table 3: Comparison of semantics-incorporated techniques on the 3DPW dataset (Von Marcard et al., 2018) based on HMR network. Building on CycleAdapt (Nam et al., 2023), we design two methods to enhance pose semantics with generated motion sequences from (Jiang et al., 2023). “downtown_rampAndStairs” and “downtown_bar” highlight different challenges. The former involves depth ambiguity when viewing people ascending stairs from behind, while the latter includes obstacles in a low-lighting bar environment. Our method consistently improves across videos while others do not.
@@
Motion discriminator (Kocabas et al., 2020) 85.7 52.9 117.9 72.2 115.0 69.1
Unpaired local poses (Jiang et al., 2023) 82.3 50.6 102.6 71.7 88.2 54.5
@@ 5.4 Analysis Experiments
[…]
87.7 53.8
79.3 49.6
78.0 48.7
76.4 47.2
Analysis on semantics-incorporated techniques. To the best of our knowledge, no work has explicitly used semantic information to improve 3D motion. Therefore, we design two other methods for semantics enhancement based on CycleAdapt (Nam et al., 2023) to compare the effectiveness with our method. Specifically, we use (Jiang et al., 2023) to obtain additional motion sequences with the same semantics as the video segment, and enforce the predicted motions to be aligned with the generated motions.
We implement the first comparison using motion discriminator (Kocabas et al., 2020). It guides motion prediction by distinguishing between predicted and generated motions with the same semantic meaning. As shown in Tab. 3, even though some cases show improvement, the overall performance is worse than CycleAdapt. One reason is the difficulty of stabilizing adversarial training; the other is that the discriminator, pre-trained to refine unrealistic motion rather than semantic differentiation.
@@
Table 5: Comparison of fill-in threshold influence on model performance (MPJPE) on the 3DPW dataset. Balancing 2D pose quantity and quality can boost model performance.
@@
Figure 4: Qualitative comparison on one video sequence from 3DPW dataset (Von Marcard et al., 2018) with text label “Walking”. We first show the necessity of 2D pose updates to preserve motion semantics. We use green and red squares to highlight the transition from walking to standing poses and demonstrate that this issue is alleviated with 2D pose updates (Ours). Next, we show that our method predicts motion more consistent with actual walking sequences compared to CycleAdapt (Nam et al., 2023).
Analysis on 2D pose update. Tab. 5 illustrates the impact of the threshold on overall performance. We find out that using 0.75 threshold adds only 36% more 2D poses, yet the high quality (PCK=0.58) helps achieve 76.4mm MPJPE. Even with a moderate quality (PCK=0.53), adequate new poses can still positively impact the model when the threshold is 0.60. Overall, the experiment demonstrates that balancing the quality and quantity of 2D poses can enhance model performance.
@@ 5.5 Qualitative Results
We perform a qualitative comparison to visually assess the effectiveness of our method. In Fig. 4, we first illustrate that without 2D pose updates, the model predictions for the truncated part revert to average poses. Specifically, in Epoch 2, frames highlighted with a green square show some walking poses. However, by epoch 4, these walking poses are back to standing poses (red square). With 2D pose updates (“Ours”), more frames accurately reflect the intended action. This result verifies the necessity of using 2D pose update to preserve motion semantics.
Furthermore, we demonstrate that CycleAdapt (Nam et al., 2023) often predicts standing poses where 2D pose estimates are missing. Our method consistently estimates motions closer to the ground truth, verifying the feasibility of using semantic information to refine predicted motions. More visualizations can be found in Supplementary Materials.
@@ 6 Conclusion and Limitations
However, our work has certain limitations: the refined motion under occlusion may not fully capture the actual pattern of the subject, with discrepancies in step frequency or duration. In addition, the reliance on MotionCLIP limits adaptability to new motion descriptions. These challenges highlight the need for better adaptation strategies and advanced motion-language models.

### v2

@@ 1 Introduction
This paper tackles the adaptation of 3D human pose estimation models to in-the-wild videos. A practical alternative to improve predictions is Test-Time Adaptation (TTA). TTA directly uses the (unlabeled) target sequences to fine-tune the pre-trained model during inference. Often, pre-trained models exhibit a misalignment between their projected keypoints and 2D image evidence (Chen et al., 2025a). Based on this result, previous TTA methods (Guan et al., 2022; Nam et al., 2023) usually penalize the distance between the 2D-projections and 2D evidence provided by either ground truth or estimated 2D poses. However, such a scheme has two main challenges (Chen et al., 2023): 1) Depth ambiguity, which arises because numerous 3D pose solutions correspond to the same 2D pose. This ambiguity may improve 2D alignment but worsen 3D accuracy and plausibility. 2) Missing 2D keypoints, e.g., under occlusions and truncations. These keypoints have no guidance for improvement.
@@ 2.1 Test-time Adaptation in 3D Pose Estimation
However, ground truth 2D poses are not always available in many real-world scenarios. Similar to our method, CycleAdapt (Nam et al., 2023) operates on estimated 2D data and introduces a cyclic framework, where the motion denoising module smooths 3D pseudo labels to mitigate the impact of noisy 2D inputs. Nonetheless, the inherent 2D-to-3D ambiguity still significantly degrades model performance, and unguided predictions often occur when 2D inputs are unreliable. Our method addresses this by leveraging explicit semantic understanding, i.e., actions, to guide the model in predicting specific motion patterns. This is the first approach to emphasize high-level semantic information to reduce the 2D-to-3D solution space in test-time 3D human pose estimation.
@@ 2.2 Human Motion Priors
Leveraging motion prior as supervision is commonly applied in video-based human pose estimation due to the lack of 3D annotations with RGB input. Trained on a large and diverse human motion database AMASS (Mahmood et al., 2019; Ionescu et al., 2013), the human prior provides a regularized motion space. Previous works mainly drag the predicted or generated motion to the regularized motion space to achieve the refinement (Kocabas et al., 2020; Shin et al., 2024; Rempe et al., 2021), infilling (Yuan et al., 2022; Zhang et al., 2021), denoising (Nam et al., 2023) goals. Specifically, VIBE (Kocabas et al., 2020) extends VAE (Pavlakos et al., 2019) and adversarial (Kanazawa et al., 2018) feasible prior from pose into motion space to improve temporal continuity of movement. GLAMR (Yuan et al., 2022) learns a VAE motion infiller to infill motion where the moving human is occluded. CycleAdapt utilizes the prior to denoise 3D meshes and improves human pose estimation during test-time with noisy 2D poses. However, those applications of human motion prior only optimize the predicted meshes of the segment to be valid motion without consideration of the motion semantics. Compared to other works, we are the first to split regularized motion space according to text labels extracted from the video segment to benefit motion prediction.
Figure 2: Overall framework: we perform model fine-tuning and 2D pose update in each epoch. Model Fine-tuning (Left): The extracted video segment and its text label will be used to update $f^{\text{HPE}}$ with 2D projection, alignment, and smooth losses. 2D Pose Update (Right): After model fine-tuning in each epoch, all images in the current test video are processed through $f^{\text{HPE}}$ to obtain predicted 2D and update the 2D keypoints using EMA. Subsequently, all video segments are checked to fill in any missing 2D keypoints by verifying conditions of motion-text alignment.
@@ 3.1 Problem Setup
3.2 Motion-language Model
A motion-language model integrates human motion and natural language modality. It utilizes pre-trained language models (Radford et al., 2018; Kenton & Toutanova, 2019), to incorporate rich semantic space for improved motion representation. This is achieved by training with motion-text data, which enables tasks like text-conditioned motion generation (Zhang et al., 2024, 2023) and motion captioning (Guo et al., 2022; Jiang et al., 2023). Our focus is on leveraging motion-language model to guide motion semantics for unlabeled test videos in TTA. We choose MotionCLIP (Tevet et al., 2022) for its text-aligned motion representation, which is optimally suited to our task.
@@ 4.1 Semantics-aware Motion Prior
More details about the VLM prompts and label verification can be found in Appendix Sec. B.
Video segment processing. After the text label for motion in each image is obtained, we aim to align the the projected motion and text features within the shared space of MotionCLIP. To process video segments for training, we divide the video into segments and only keep those where all images share the same text label. This is to avoid misalignment caused by mixed motions as much as possible. To further improve alignment, we adopt weighted sampling and prioritize those segments whose motion predictions deviate from the true semantics. Specifically, the sampling weight is defined as 1 minus cosine similarity between predicted motion and its text label in MotionCLIP.
@@ 4.3 Overall Training
${\mathcal{L}}_{\textnormal{smooth}}=||\bm{\hat{\theta}}-\bm{\bar{\theta}}||_{1}+||\bm{\hat{\beta}}-\bm{\bar{\beta}}||_{1},~{}$ (9)
where the $\bm{\bar{\theta}}$ comes from the motion denoise module, and $\bm{\bar{\beta}}$ is the averaged shape estimation with a sliding window operation. Together with motion-text feature alignment loss, 2D projection loss, and loss weight $\lambda$, our training objective in each epoch can be summarized as:
@@ 5 Experiments
Table 2: Test-time adaptation comparison results on EgoBody (Zhang et al., 2022) and 3DPW (Von Marcard et al., 2018) based on HMR2.0 network (Goel et al., 2023). Our method achieves better performance than CycleAdapt (Nam et al., 2023).
@@ 5.1 Setup
Datasets. We follow the adaptation tasks from previous work (Zhang et al., 2020; Nam et al., 2023), using Human3.6M (Ionescu et al., 2013) as the labeled training dataset and 3DPW (Von Marcard et al., 2018) and 3DHP (Mehta et al., 2017) as the unlabeled test datasets. Human3.6M is a widely used indoor dataset comprising 3.6 million images annotated with 2D and 3D labels. Despite its extensive pose diversity, the HMR network (Kanazawa et al., 2018) pre-trained on this dataset often struggles to generalize to in-the-wild settings due to its limited variation in appearance and environmental factors. The 3DPW test set comprises 37 videos and presents a significant challenge as it is an outdoor dataset featuring diverse scenarios, including subject occlusion and varied environments. The 3DHP test set includes 6 videos with both indoor and outdoor scenarios, and it presents an additional challenge by featuring some poses not seen in Human3.6M. Furthermore, we validate our method on a egocentric dataset EgoBody (Zhang et al., 2022), which presents severe body truncation with missing 2D evidence.
@@ 5.2 Implementation Details
At the start of test-time adaptation for each test video, the model parameters are initialized with the pre-trained values, following (Nam et al., 2023). We employed the Adam optimizer (Kingma & Ba, 2014) with parameters set to $beta1$ = 0.5, $beta2$ = 0.9, and a learning rate of 5.0e-5. A cosine scheduler is used with a minimum learning rate of 1.0e-6. The input images are resized to 224$\times$224, and the frame number of each video segment is 60. We use a batch size of 4 and the total training epoch is 6. The hyperparameters are $\lambda_{1}=0.1$, $\lambda_{2}=0.2$, $\sigma=0.75$, $\alpha=0.9$. We use Openpose (Cao et al., 2017) to provide 2D poses with a confidence threshold of 0.3 (Gu et al., 2024). The comparison methods include state-of-the-art test-time adaptation methods: BOA (Guan et al., 2021), DynaBOA (Guan et al., 2022), DAPA (Weng et al., 2022) and CycleAdapt (Nam et al., 2023).
[…]
87.7 53.8
79.3 49.6
78.0 48.7
76.4 47.2
Table 4: Analysis of semantics-incorporated strategies on the 3DPW dataset (Von Marcard et al., 2018) based on HMR network. Building on CycleAdapt (Nam et al., 2023), we design two methods to enhance pose semantics with generated motion sequences from (Jiang et al., 2023). “downtown_rampAndStairs” and “downtown_bar” highlight different challenges. The former involves depth ambiguity when viewing people ascending stairs from behind, while the latter includes obstacles in a low-lighting bar environment. Our method consistently improves across videos while others do not.
@@
+Motion discriminator (Kocabas et al., 2020) 85.7 52.9 117.9 72.2 115.0 69.1
+Unpaired local poses (Jiang et al., 2023) 82.3 50.6 102.6 71.7 88.2 54.5
@@ 5.4 Analysis Experiments
Semantics-incorporated strategies analysis. To the best of our knowledge, no existing works explicitly use semantic information to improve 3D motion in TTA. Therefore, we verify our method through comparison to the most relevant components from the literature, as shown in Tab. 4. Specifically, we use CycleAdapt (Nam et al., 2023) as a baseline, which lacks explicit semantic guidance. We then evaluate two alternative strategies for incorporating semantics, replacing our proposed motion-text alignment with each: motion discriminator (Kocabas et al., 2020) (second row) and unpaired local poses (third row). Through adversarial loss or L1 minimization, the predicted motion is enforced to align with the generated motion from (Jiang et al., 2023) by sharing the same semantics as the video segment.
We implement the first comparison using motion discriminator. It guides motion prediction by distinguishing between predicted and generated motions of the same semantic meaning. As shown in Tab. 4, despite improvement in some cases, the overall performance is worse than CycleAdapt. One reason is the difficulty of stabilizing adversarial training; the other is that the discriminator, pre-trained to refine unrealistic motion rather than semantic differentiation.
Table 5: Runtime analysis (ms) on the 3DPW dataset. Our method introduces a modest runtime increase compared to CycleAdapt, but is still much faster than other TTA methods.
Method BOA DynaBOA DAPA CycleAdapt Ours
Runtime 840.3 1162.8 431.0 74.1 107.6
Figure 4: The relationship between the fill-in threshold, number of keypoints, and 2D keypoint PCK. Increasing the threshold leads to a decrease in the ratio of filled-in keypoints, but an improvement in their quality, as indicated by a higher PCK.
Runtime analysis. We follow the same runtime evaluation setup as CycleAdapt (Nam et al., 2023) to ensure fair comparison. As shown in Tab. 5, methods like BOA (Guan et al., 2021), DynaBOA (Guan et al., 2022), and DAPA (Weng et al., 2022) have high per-frame runtime due to complex optimization or extra synthesis rendering. Compared to CycleAdapt, our method introduces an additional 33.5ms runtime per frame, 85.4% of which comes from the VLM, and 14.6% from our proposed framework components. However, it is still much faster than other TTA methods.
2D fill-in analysis. We try different fill-in thresholds in Eq. 8 (from 0.55 to 0.80) to evaluate the quality and quantity of 2D keypoints. Fig. 4 shows as we increase the fill-in threshold, fewer new 2D keypoints are added (blue bars), but their PCK improves (green bars). Increasing the threshold helps use reliable but fewer predicted 2D keypoints in training, which have high feature similarity with their text labels.
Tab. 6 illustrates the overall performances on different thresholds. We find out that using 0.75 threshold adds only 36% more 2D keypoints, yet the high quality (PCK=0.58) helps achieve 76.4mm MPJPE. Even with a moderate quality (PCK=0.53), adequate new keypoints can still positively impact the model when the threshold is 0.60. Overall, the experiment demonstrates that balancing the quality and quantity of 2D keypoints can enhance model performance.
Table 6: Comparison of fill-in threshold influence on model performance (MPJPE) on the 3DPW dataset. Balancing 2D keypoint quantity and quality can boost model performance.
@@
Table 7: Comparison on VLM failure cases: initial performance without adaptation, performance after adaptation with incorrect VLM label and correct (oracle) label, respectively. Adaptation with incorrect VLM label can still reduce MPJPE; accurate label yield greater improvement.
Method MPJPE
Initial 213.8
Incorrect VLM label 92.2
Correct (oracle) label 83.7
5.5 VLM Failure Cases
We manually examined 5,000 frames from the 3DPW dataset, where 96.4% is accurate, as the actions are relatively simple and VLM is reliable. For the remaining 3.6%, the VLM hallucinates actions based on background, e.g. a person picking up foil is mistaken as “golfing” as the foil resembles a golf club, or “eating something using both hands” and “cast fishing pole” in scenes with fruit stands and rivers, respectively. The visualization of failure cases can be found in Appendix Fig. 11.
Despite these VLM failures, our method maintains robustness. As our semantics-aware alignment is a regularizer, its impact is limited if it conflicts with 2D projection loss. Tab. 7 shows that the adaptation with the wrong label still reduces the MPJPE, but not as much as if the correct (oracle) action label were applied.
Figure 5: Qualitative comparison on one video sequence from 3DPW dataset (Von Marcard et al., 2018) with text label “Walking”. From epoch 2 to epoch 4, results show that motion-text alignment alone is insufficient to preserve motion semantics with missing 2D data. We use green and red squares to highlight the transition from walking to standing poses. Incorporating 2D pose updates (Ours) helps mitigate this problem. Compared to CycleAdapt (Nam et al., 2023), which shows static standing foot under image truncation, ours captures the natural alternation of foot and knee bending during walking, aligning more closely with the actual walking sequences.
@@ 5.6 Qualitative Results
We perform a qualitative comparison to visually assess the effectiveness of our method. In Fig. 5, we first illustrate that without 2D pose updates, the model predictions for the truncated part revert to average poses. Specifically, in epoch 2, frames highlighted with a green square show some walking poses. However, by epoch 4, these walking poses are back to standing poses (red squares). With 2D pose updates (“Ours”), more frames accurately reflect the intended action. This result verifies the necessity of using 2D pose update to preserve motion semantics.
Furthermore, we demonstrate that CycleAdapt (Nam et al., 2023) often predicts standing poses where 2D pose estimates are missing. Our method consistently estimates motions closer to the ground truth, verifying the feasibility of using semantic information to refine predicted motions. More visualizations can be found in Appendix Sec. C.
@@ 6 Conclusion and Limitations
However, our work has certain limitations: the refined motion under occlusion may not fully capture the actual pattern of the subject, with discrepancies in step frequency or duration; ambiguities like motion blur (Chen et al., 2025b) remains an open direction; and reliance on MotionCLIP limits adaptability to new motion descriptions. Addressing these challenges calls for improved adaptation strategies and more advanced motion-language models.
Acknowledgments
This research was funded in part by a Google South Asia & Southeast Asia Research Award given to A. Yao.